OpenStudy (anonymous):

Find the eqeuation of the parabola that passes theough (3,0)(-5,0) and (5,-50)

OpenStudy (whpalmer4):

If the parabola passes through (3,0) and (-5,0), that means the equation of the parabola has roots at x = 3 and x = -5. We can write the equation as \[P(x) = k(x-3)(x+5)\] where \(k\) is a scaling factor to get the parabola to hit our target point at (5,-50). Multiply out the polynomial, and find the value of \(k\) such that \(P(5) = -50\).

OpenStudy (whpalmer4):

Actually, no need to multiply it out before finding \(k\), probably easier if you don't!

OpenStudy (anonymous):

I'm a little confused how do you find k

OpenStudy (whpalmer4):

\[P(5) = -50 = k(x-3)(x+5) = k(5-3)(5+5) = 20k\]\[k=?\]

OpenStudy (anonymous):

20?

OpenStudy (whpalmer4):

Might be confusing to read with so many versions on the same line: \[-50 = 20k\] Solve for \(k\)

OpenStudy (anonymous):

k=-2.5

OpenStudy (whpalmer4):

Exactly! So what is your final equation?

OpenStudy (whpalmer4):

(The value of k also controls whether the parabola opens up or down, of course)

OpenStudy (anonymous):

p(x)=-2.5(x-3)(x+5) ?

OpenStudy (anonymous):

Is that correct?

OpenStudy (whpalmer4):

Yes, or (-5/2)(x^2+2x-15) or whatever. Here's a plot of the final result:

OpenStudy (whpalmer4):

If you squint just right you can see that the curve goes through (5,-50) and crosses the x-axis at x = -5 and x = 3.

OpenStudy (anonymous):

ok thanks

OpenStudy (whpalmer4):

Does that make sense to you?